# REPS: Recycled Entropy Packet Spraying for Adaptive Load Balancing and Failure Mitigation

### v3

@@ Appendix A Freezing Mode in REPS
In our paper, we focused on scenarios where packet trimming was not supported. However, REPS performs optimally when packet trimming is available, benefiting from both an enhanced loss detection algorithm and a more responsive CC loop.
Regardless of the employed strategy, REPS maintains high performance even if it inadvertently enters freezing mode without an actual network failure. This is because entering freezing mode effectively reduces the EVS size of REPS. As demonstrated in Section 4.5.1, REPS remains effective with as few as 32 EVs. For instance, we tested REPS with the 16 MiB tornado workload, running it first under normal conditions and then with forced freezing mode activated after $150\;\mu\text{s}$. The results showed only a 1% increase in completion time for REPS with forced freezing mode, with the 99th percentile RTT remaining very close, as illustrated in Figure 15.
As an extension for REPS, probing can be incorporated to make the failure detection more precise, but we decided to not add this as part of REPS for now in order to keep things simple. Moreover, in this paper we do not discuss fast loss recovery mechanisms as they are anyway orthogonal to REPS behaviour and could be used to further improve its performance.
Figure 15. Studying REPS while forcing freezing mode without failures.
@@ Appendix B Entropy Values Set Analysis
We first show theoretically that the EVS size is very important for OPS and later that REPS is much more flexible and can work well even with small EVS.
@@
Appendix C REPS Single Switch Convergence Proof
@@ D.1. ACK Coalescing Theoretical Modeling
To assess REPS performance under different ACK coalescing ratios, we validate it using the theoretical model from Section 5.1.
@@ D.2. Different Tiers
We aim to verify that REPS performs effectively with fat-tree topologies that have three tiers. This scenario poses a slightly greater challenge for REPS, as a single EV must manage two hops. Nonetheless, there is no intrinsic reason why REPS should not perform well in such a topology.
To validate this, we execute the synthetic benchmark using the symmetric topology described in Section 4.3.1, but with three tiers instead of two. The results, shown in Figure 18, indicate that REPS performs comparably to the two-tier topology.
Figure 18. Studying REPS with a 3 tiers fat tree topology.
@@ D.3. Incremental Failures
To further demonstrate the resilience of REPS under failures, we conducted an experiment where we incrementally failed all but one of a switch’s uplinks at 200 µs intervals. Figure 19 presents the permutation from the perspective of the failing switch, where three of the four uplinks were permanently disabled in a staggered manner. As expected, REPS enters freezing mode immediately after the first failure, ensuring that failing output ports are avoided. Notably, small utilization spikes are observed on the failing links when REPS exits freezing mode to verify if the failure has been resolved. Since the failures are permanent for the duration of the experiment, REPS promptly re-enters freezing mode after detecting unresolved issues.
In this drastic scenario, OPS performs 40× worse than REPS, primarily due to its inability to avoid broken links, resulting in numerous retransmissions and reduced congestion windows.
Figure 19. REPS vs. OPS in a 32 MiB permutation with incremental persistent failures.
@@ Appendix E Additional Data
The datacenter traces used throughout this paper have been previously used in a number of similar works (Vanini et al., 2017; Alizadeh et al., 2010a). In particular we use traces provided by Alibaba and Facebook and a series of traces used for web search in production clusters. The CDF distribution for such traces can be seen in Figure 20. For most of the paper we focus exclusively on the WebSearch traces.

### v4

@@ Appendix A Freezing Mode in ARCANE
In our paper, we focused on scenarios where packet trimming was not supported. However, ARCANE performs optimally when packet trimming is available, benefiting from both an enhanced loss detection algorithm and a more responsive CC loop.
Regardless of the employed strategy, ARCANE maintains high performance even if it inadvertently enters freezing mode without an actual network failure. This is because entering freezing mode effectively reduces the EVS size of ARCANE. As demonstrated in Section 4.5.2, ARCANE remains effective with as few as 32 EVs. For instance, we tested ARCANE with the 16 MiB tornado workload, running it first under normal conditions and then with forced freezing mode activated after $150\;\mu\text{s}$. The results showed only a 1% increase in completion time for ARCANE with forced freezing mode, with the 99th percentile RTT remaining very close, as illustrated in Figure 15.
As an extension of ARCANE, probing can be incorporated to enhance failure detection precision. However, we chose not to integrate this into ARCANE at this stage to maintain clarity and focus on the core framework. We also do not discuss fast loss-recovery mechanisms as they are anyway orthogonal to ARCANE behaviour and could be used to further improve its performance.
Figure 15. Studying ARCANE while forcing freezing mode without failures.
@@ Appendix B Entropy Values Set Analysis
We first show theoretically that the EVS size is very important for OPS and later that ARCANE is much more flexible and can work well even with small EVS.
@@
Appendix C ARCANE Single Switch Convergence Proof
@@ D.1. ACK Coalescing Theoretical Modeling
To assess ARCANE performance under different ACK coalescing ratios, we validate it using the theoretical model from Section 5.1.
@@ D.2. Different Tiers
We aim to verify that ARCANE performs effectively with fat-tree topologies that have three tiers. This scenario poses a slightly greater challenge for ARCANE, as a single EV must manage two hops. Nonetheless, there is no intrinsic reason why ARCANE should not perform well in such a topology.
To validate this, we execute the synthetic benchmark using the symmetric topology described in Section 4.3.1, but with three tiers instead of two. The results, shown in Figure 18, indicate that ARCANE performs comparably to the two-tier topology.
Figure 18. Studying ARCANE with a 3 tiers fat tree topology.
@@ D.3. Incremental Failures
To further demonstrate the resilience of ARCANE under failures, we conducted an experiment where we incrementally failed all but one of a switch’s uplinks at 200 µs intervals. Figure 19 presents the permutation from the perspective of the failing switch, where three of the four uplinks were permanently disabled in a staggered manner. As expected, ARCANE enters freezing mode immediately after the first failure, ensuring that failing output ports are avoided. Notably, small utilization spikes are observed on the failing links when ARCANE exits freezing mode to verify if the failure has been resolved. Since the failures are permanent for the duration of the experiment, ARCANE promptly re-enters freezing mode after detecting unresolved issues.
In this drastic scenario, OPS performs 40× worse than ARCANE, primarily due to its inability to avoid broken links, resulting in numerous retransmissions and reduced congestion windows.
Figure 19. ARCANE vs. OPS in a 32 MiB permutation with incremental persistent failures.
@@ Appendix E Additional Data
The datacenter traces used throughout this paper have been previously used in a number of similar works (Vanini et al., 2017; Alizadeh et al., 2010). In particular we use traces provided by Facebook and a series of traces used for web search in production clusters. The CDF distribution for such traces can be seen in Figure 20. For most of the paper we focus exclusively on the WebSearch traces.